# Dynamic Prediction for Hospital Readmission in Patients with Chronic Heart Failure

### v1

@@ Abstract
Hospital readmission among patients with chronic heart failure (HF) is a major clinical and economic burden. Dynamic prediction models that leverage longitudinal biomarkers may improve risk stratification over traditional static models. This study aims to develop and validate a joint model (JM) using longitudinal N-terminal pro-B-type natriuretic peptide (NT-proBNP) measurements to predict the risk of rehospitalization or death in HF patients. We analyzed real-world data from the TriNetX database, including patients with an incident HF diagnosis between 2016 and 2022. The final selected cohort included 1,804 patients. A Bayesian joint modeling framework was developed to link patient-specific NT-proBNP trajectories to the risk of a composite endpoint (HF rehospitalization or all-cause mortality) within a 180-day window following hospital discharge. The model’s performance was evaluated using 5-fold cross-validation and assessed with the Integrated Brier Score (IBS) and Integrated Calibration Index (ICI). The joint model demonstrated a strong predictive advantage over a benchmark static model, particularly when making updated predictions at later time points (180-360 days). A joint model trained on patients with more frequent NT-proBNP measurements achieved the highest accuracy. The main joint model showed excellent calibration, suggesting its risk estimates are reliable. These findings suggest that modeling the full trajectory of NT-proBNP with a joint modeling framework enables more accurate and dynamic risk assessment compared to static, single-timepoint methods. This approach supports the development of adaptive clinical decision-support tools for personalized HF management.
Introduction
@@
Chronic heart failure (HF) represents a significant challenge in contemporary healthcare systems due to its high prevalence, disease burden, and frequent hospital readmissions. HF affects approximately 2% of the adult population worldwide, with prevalence rising to nearly 10% among individuals over 75 years old [1, 2]. The condition is characterized by impaired cardiac function leading to symptoms such as dyspnea, fatigue, and fluid retention [3]. Despite advances in therapy, HF remains associated with substantial morbidity and mortality. Notably, up to 50% of patients experience hospital readmission within 180 days following discharge [4], and the early post-discharge period represents a clinically vulnerable phase marked by disproportionately high event rates [5]. Accurate prediction of hospital readmission is crucial for improving clinical outcomes and reducing healthcare costs, highlighting the need for advanced predictive methodologies.
@@
Dynamic prediction through joint modeling (JM) has gained considerable attention as a powerful statistical framework that simultaneously analyzes longitudinal biomarkers and time-to-event data. A shared random effects joint model links a survival process with a longitudinal process via common latent variables, typically random effects, to capture individual-specific trajectories and their influence on event risk. Early foundational work on joint models focused on a single longitudinal outcome modeled via linear mixed-effects models [6, 7]. Building on this work, subsequent research expanded the methodology to incorporate Bayesian estimation techniques, which allow flexible inference and dynamic prediction [8, 9, 10, 11, 12, 13]. These frameworks have been further extended to handle more complex data structures, including competing risks [14, 15], and multiple longitudinal biomarkers [16, 17]. The availability of open-source software has been pivotal in applying joint models in practice. Notably, the JM package [18] in R was among the first tools for fitting joint models, while the more recent JMbayes2 package [19] offers Bayesian joint models with enhanced flexibility.
@@
In the context of heart failure, dynamic prediction models utilizing JM frameworks have shown great promise by leveraging longitudinal biomarker data to update risk assessments over time. Central to these models is N-terminal pro-B-type natriuretic peptide (NT-proBNP), a cornerstone biomarker released from cardiac myocytes in response to increased wall stress [20, 21]. Higher levels of NT-proBNP reflect ongoing myocardial strain and are strongly correlated with disease severity and mortality risk [20, 21]. While NT-proBNP is widely used clinically, most prognostic models rely on a single baseline measurement, which may overlook the evolving risk as disease progresses and treatments change. Indeed, the prognostic importance of changes in NT-proBNP over time, not just its baseline value, was demonstrated in major clinical trials and subsequent meta-analyses [22, 23]. This finding provides a strong rationale for more sophisticated analytical methods. A few studies have applied joint models to HF datasets, integrating longitudinal NT-proBNP measurements to generate dynamically updated risk predictions for adverse clinical events [24, 25, 26, 27]. Joint models offer a powerful approach to capture these dynamic risk patterns and therefore improve individualized prognosis. Beyond HF, joint models have also been effectively applied to other cardiovascular conditions. For example, JM-based dynamic prediction has been extended to congenital heart disease data, leveraging NT-proBNP to improve prognostic accuracy [28]. More recently, the field has advanced through the incorporation of competing risks into JM dynamic predictions for HF, further enhancing model applicability [29]. These advances underscore the potential of dynamic prediction models to offer individualized, time-adaptive risk predictions that can support tailored clinical decision-making.
@@
Our study builds on the existing literature by developing a robust JM framework to predict hospital readmissions in patients with chronic heart failure. Specifically, we investigate both individual and population-level longitudinal dynamics of NT-proBNP in relation to HF-related rehospitalization, evaluate whether patient-specific NT-proBNP trajectories from hospital discharge to readmission can dynamically predict clinical outcomes, and identify the most informative functional form linking biomarker evolution to event risk. Indeed, whether absolute levels or relative changes in NT-proBNP serve as more important drivers of risk is an aspect that remains unresolved in previous studies [24, 26]. By leveraging dynamic prediction, we intend to better characterize patient-specific evolution and empower clinicians with adaptive tools for optimizing patient management.
Methods
Data preprocessing and filtering
@@ Methods
Endpoint definition and estimand of interest
@@
Longitudinal Biomarker: NT-proBNP
The primary longitudinal biomarker analyzed in this study is NT-proBNP, a well-established indicator strongly associated with heart failure outcomes [24, 25, 26, 29]. To be included in the analysis, patients were required to have at least one NT-proBNP measurement within 30 days before their first HF hospitalization ($t_{0}$) and a minimum of two measurements after $t_{0}$ but before the occurrence of the event or the last follow-up visit. These conditions ensure sufficient longitudinal data to capture individual biomarker trajectories. In cases where multiple NT-proBNP measurements were recorded on the same date for a patient, the average value is used to represent that day. Furthermore, values of NT-proBNP exceeding 70,000 pg/mL were excluded from the analysis based on clinical expertise, as such extreme levels may indicate measurement anomalies. After applying these NT-proBNP related inclusion criteria (following prior exclusions for data inconsistencies and follow-up loss), approximately 8% of the initial cohort remained eligible for our study.
Baseline characteristics
@@
Table 1: Summary of baseline demographic and clinical characteristics for the study cohort (n=1,804). Continuous variables are presented as mean (standard deviation), while categorical variables are shown as counts (percentages). All features were measured within 30 days prior to the landmark time (date of first HF hospitalization discharge).
@@
Statistical methods
Joint modeling of longitudinal and survival data
The JM framework offers a powerful approach to simultaneously analyze the evolution of a longitudinal biomarker and its direct impact on the time to a clinical event, while properly accounting for measurement error and individual-level variability. Let $T_{i}$ denote the observed event time for the $i$-th individual, with $i=1,\dots,n$, defined as the minimum between the true event time $T_{i}^{*}$ and the censoring time $C_{i}$, i.e., $T_{i}=\min(T_{i}^{*},C_{i})$. The event indicator is defined as $\delta_{i}=I(T_{i}^{*}\leq C_{i})$. The longitudinal biomarker measurement (NT-proBNP) for individual $i$ at time $t$ (in days) is denoted by $y_{i}(t)$. The observed longitudinal data for individual $i$ consists of measurements $y_{i}=\{y_{i}(t_{ij})\}_{j=1}^{n_{i}}$ collected at times $t_{i}=\{t_{ij}\}_{j=1}^{n_{i}}$, where $n_{i}$ is the number of observations for that individual. Hence, the complete observed data for subject $i$ is $(T_{i},\delta_{i},y_{i},t_{i})$. We denote by $m_{i}(t)$ the true and unobserved value of the longitudinal outcome at time $t$. To assess its effect on event risk, we estimate $m_{i}(t)$ via a mixed-effects model, which combines fixed effects to capture population-level trends and random effects to characterize individual variation, a well-established approach in longitudinal data analysis [30, 31]. Specifically, we incorporate quadratic splines of time to flexibly model subject-specific trajectories:
@@
Crucially, the hazard depends on the true unobserved longitudinal process $m_{i}(t)$ rather than the noisy observations $y_{i}(t)$. This connection is established through the random effects $b_{i}$, which serve as the bridge between the longitudinal and survival submodels, capturing how individual-specific biomarker trajectories relate to their risk of experiencing the event. For the baseline hazard function, several specifications are possible. In this study, $h_{0}(t)$ is modeled parametrically using the Weibull distribution [32]. Alternatively, one could use other distributions or leave $h_{0}(t)$ unspecified, as in the Cox proportional hazards regression.
A classical approach to estimating joint model parameters relies on maximum likelihood, where the log-likelihood is derived from the joint distribution of the longitudinal and time-to-event outcomes. Obtaining the marginal likelihood requires approximation due to intractable integrals. Methods like Gaussian quadrature or Monte Carlo integration are commonly used [33, 34]. In this work, we instead adopt a Bayesian approach [8, 9], where inference is based on the posterior distribution of the model parameters and latent variables. Specifically, we use Markov Chain Monte Carlo (MCMC) methods to estimate the parameters of the proposed joint model. The joint likelihood is formulated under the assumption that the longitudinal and survival processes are conditionally independent given the subject-specific random effects.
@@
Model specification
The longitudinal submodel, describing the evolution of log-transformed NT-proBNP values over time, was implemented via the nlme[35] package. To flexibly capture the non-linear subject-specific trajectories and variability over time (see Figure 1F), we used quadratic splines of time for both fixed and random effects. We did not include additional covariates in the longitudinal model, in order to focus on the temporal dynamics of NT-proBNP and its association with event risk, and to avoid overfitting given the limited number of measurements for many patients. The survival submodel includes all pre-selected baseline covariates. We assumed a parametric Weibull baseline hazard to enable fully parametric estimation of the time-to-event distribution, implemented using the survival package [36].
@@
For benchmarking purposes, we also fitted a standard survival Weibull model using all baseline covariates and the most recent NT-proBNP measurement as an additional covariate, consistent with the approach in prior work [26], and refer to this model as the benchmark model. This comparison is intended to determine whether modeling the entire biomarker trajectory dynamically offers advantages over relying on a single static measurement. All models were estimated using the JMbayes2[19] package in R, which employs a full Bayesian framework with MCMC algorithms. Posterior inference is obtained via a combination of Gibbs sampling and Metropolis-Hastings algorithm. We used the default priors in the package JMbayes2, which primarily consist of weakly informative normal priors for the regression coefficients and half-t or gamma priors for the variance and standard deviation components.
Results
Exploratory data analysis
@@ Results
The NT-proBNP measurements in this cohort span a wide range, with a median concentration of 1757 pg/mL (IQR: 578 - 4637.5 pg/mL). In our analysis NT-proBNP values are log-transformed to ensure positivity and to better approximate a symmetric distribution, following common practice in the literature [25, 26]. Figure 1C displays the distribution of the log-transformed NT-proBNP measurements collected from any patient, totaling 9,594 values. The time distribution of these values is shown in Figure 1D, which illustrates that the time measurements of NT-proBNP are concentrated in the first part of the follow-up period, with almost 70% of the values recorded within the first 180 days after $t_{0}$. Recall that biomarker measurements taken within 30 days prior to $t_{0}$ and at $t_{0}$ are included in the analysis; in fact, approximately 31% of all measurements fall within this pre-baseline window. The average time of the first NT-proBNP measurement per patient is 9 days before $t_{0}$, while the average time of the last measurement is 367 days after $t_{0}$. The distribution of the number of NT-proBNP measurements per patient (Figure 1E) indicates that the largest proportion of patients (33%) have exactly three recorded values, which is the minimum required by our inclusion criteria. The limited number of longitudinal measurements represents a key challenge of the analysis, as it constrains our ability to robustly characterize individual biomarker trajectories over time. Individual longitudinal trajectories, shown in Figure 1F further demonstrate substantial heterogeneity in NT-proBNP patterns.
Model fit and baseline risk associations

### v2

@@ Abstract
Background: Hospital readmission among patients with chronic heart failure (HF) is a major clinical and economic burden. Dynamic prediction models that leverage longitudinal biomarkers may improve risk stratification over traditional static models. This study aims to develop and validate a joint model using longitudinal N-terminal pro-B-type natriuretic peptide (NT-proBNP) measurements to predict the risk of rehospitalization or death in HF patients.
Methods: We analyzed real-world data from the TriNetX database, including patients with an incident HF diagnosis between 2016 and 2022. The final selected cohort included 1,804 patients. A Bayesian joint modeling framework was developed to link patient-specific NT-proBNP trajectories to the risk of a composite endpoint (HF rehospitalization or all-cause mortality) within a 180-day window following hospital discharge. The model’s performance was evaluated using 5-fold cross-validation and assessed with the Integrated Brier Score and Integrated Calibration Index.
Results: The joint model demonstrated a strong predictive advantage over a benchmark static model, particularly when making updated predictions at later time points (180-360 days). A joint model trained on patients with more frequent NT-proBNP measurements achieved the highest accuracy. The main joint model showed excellent calibration, suggesting its risk estimates are reliable.
Conclusion: Our findings suggest that modeling the full trajectory of NT-proBNP with a joint modeling framework enables more accurate and dynamic risk assessment compared to static, single-timepoint methods. This approach supports the development of adaptive clinical decision-support tools for personalized HF management.
keywords:
Bayesian Joint Model, Dynamic Prediction, Heart Failure, Hospital Readmission, Individualized Risk Prediction, NT-proBNP, Real-world Data, TriNetX
1 Background
@@
Chronic heart failure (HF) represents a significant challenge in contemporary healthcare systems due to its high prevalence, disease burden, and frequent hospital readmissions. HF affects approximately 2% of the adult population worldwide, with prevalence rising to nearly 10% among individuals over 75 years old [bui2011epidemiology, metra2017heart]. The condition is characterized by impaired cardiac function leading to symptoms such as dyspnea, fatigue, and fluid retention [ponikowski20162016]. Despite advances in therapy, HF remains associated with substantial morbidity and mortality. Notably, up to 50% of patients experience hospital readmission within 180 days following discharge [desai2012rehospitalization], and the early post-discharge period represents a clinically vulnerable phase marked by disproportionately high event rates [greene2015vulnerable]. Accurate prediction of hospital readmission is crucial for improving clinical outcomes and reducing healthcare costs, highlighting the need for advanced predictive methodologies.
@@
Dynamic prediction through joint modeling (JM) has gained considerable attention as a powerful statistical framework that simultaneously analyzes longitudinal biomarkers and time-to-event data. A shared random effects joint model links a survival process with a longitudinal process via common latent variables, typically random effects, to capture individual-specific trajectories and their influence on event risk. Early foundational work on joint models focused on a single longitudinal outcome modeled via linear mixed-effects models [Wulfsohn1997, hogan1997]. Building on this work, subsequent research expanded the methodology to incorporate Bayesian estimation techniques, which allow flexible inference and dynamic prediction [rizopoulos2011dynamic, rizopoulos2012joint, barrett2017dynamic, andrinopoulou2018improved, ferrer2019individual, andrinopoulou2021reflection]. These frameworks have been further extended to handle more complex data structures, including competing risks [andrinopoulou2014, andrinopoulou2017], and multiple longitudinal biomarkers [rizopoulos2011bayesian, mauff2020joint]. The availability of open-source software has been pivotal in applying joint models in practice. Notably, the JM package [rizopoulos2010jm] in R was among the first tools for fitting joint models, while the more recent JMbayes2 package [rizopoulos2025jmbayes2] offers Bayesian joint models with enhanced flexibility.
@@
In the context of heart failure, dynamic prediction models utilizing JM frameworks have shown great promise by leveraging longitudinal biomarker data to update risk assessments over time. Central to these models is N-terminal pro-B-type natriuretic peptide (NT-proBNP), a cornerstone biomarker released from cardiac myocytes in response to increased wall stress [maisel2002rapid, felker2006natriuretic]. Higher levels of NT-proBNP reflect ongoing myocardial strain and are strongly correlated with disease severity and mortality risk [maisel2002rapid, felker2006natriuretic]. While NT-proBNP is widely used clinically, most prognostic models rely on a single baseline measurement, which may overlook the evolving risk as disease progresses and treatments change. Indeed, the prognostic importance of changes in NT-proBNP over time, not just its baseline value, was demonstrated in major clinical trials and subsequent meta-analyses [masson2008prognostic, zile2016prognostic]. This finding provides a strong rationale for more sophisticated analytical methods. A few studies have applied joint models to HF datasets, integrating longitudinal NT-proBNP measurements to generate dynamically updated risk predictions for adverse clinical events [brankovic2018, vanBoven2018, zhang2018, fuery2024prognostic]. Joint models offer a powerful approach to capture these dynamic risk patterns and therefore improve individualized prognosis. Beyond HF, joint models have also been effectively applied to other cardiovascular conditions. For example, JM-based dynamic prediction has been extended to congenital heart disease data, leveraging NT-proBNP to improve prognostic accuracy [baggen2018]. More recently, the field has advanced through the incorporation of competing risks into JM dynamic predictions for HF, further enhancing model applicability [petersen2025]. These advances underscore the potential of dynamic prediction models to offer individualized, time-adaptive risk predictions that can support tailored clinical decision-making.
@@
Our study builds on the existing literature by developing a robust JM framework to predict hospital readmissions in patients with chronic heart failure. Specifically, we investigate both individual and population-level longitudinal dynamics of NT-proBNP in relation to HF-related rehospitalization, evaluate whether patient-specific NT-proBNP trajectories from hospital discharge to readmission can dynamically predict clinical outcomes, and identify the most informative functional form linking biomarker evolution to event risk. Indeed, whether absolute levels or relative changes in NT-proBNP serve as more important drivers of risk is an aspect that remains unresolved in previous studies [brankovic2018, zhang2018]. By leveraging dynamic prediction, we intend to better characterize patient-specific evolution and empower clinicians with adaptive tools for optimizing patient management.
2 Methods
2.1 Data preprocessing and filtering
@@
2.1.1 Endpoint definition and estimand of interest
@@
2.1.2 Longitudinal Biomarker: NT-proBNP
The primary longitudinal biomarker analyzed in this study is NT-proBNP, a well-established indicator strongly associated with heart failure outcomes [brankovic2018, vanBoven2018, zhang2018, petersen2025]. To be included in the analysis, patients were required to have at least one NT-proBNP measurement within 30 days before their first HF hospitalization ($t_{0}$) and a minimum of two measurements after $t_{0}$ but before the occurrence of the event or the last follow-up visit. These conditions ensure sufficient longitudinal data to capture individual biomarker trajectories. In cases where multiple NT-proBNP measurements were recorded on the same date for a patient, the average value is used to represent that day. Furthermore, values of NT-proBNP exceeding 70,000 pg/mL were excluded from the analysis based on clinical expertise, as such extreme levels may indicate measurement anomalies. After applying these NT-proBNP related inclusion criteria (following prior exclusions for data inconsistencies and follow-up loss), approximately 8% of the initial cohort remained eligible for our study.
2.1.3 Baseline characteristics
@@
Table 1: Summary of baseline demographic and clinical characteristics for the study cohort (n=1,804).00footnotetext: Continuous variables are presented as mean (standard deviation), while categorical variables are shown as counts (percentages). All features were measured within 30 days prior to the landmark time (date of first HF hospitalization discharge).
@@
2.2 Statistical methods
2.2.1 Joint modeling of longitudinal and survival data
The JM framework offers a powerful approach to simultaneously analyze the evolution of a longitudinal biomarker and its direct impact on the time to a clinical event, while properly accounting for measurement error and individual-level variability. Let $T_{i}$ denote the observed event time for the $i$-th individual, with $i=1,\dots,n$, defined as the minimum between the true event time $T_{i}^{*}$ and the censoring time $C_{i}$, i.e., $T_{i}=\min(T_{i}^{*},C_{i})$. The event indicator is defined as $\delta_{i}=I(T_{i}^{*}\leq C_{i})$. The longitudinal biomarker measurement (NT-proBNP) for individual $i$ at time $t$ (in days) is denoted by $y_{i}(t)$. The observed longitudinal data for individual $i$ consists of measurements $y_{i}=\{y_{i}(t_{ij})\}_{j=1}^{n_{i}}$ collected at times $t_{i}=\{t_{ij}\}_{j=1}^{n_{i}}$, where $n_{i}$ is the number of observations for that individual. Hence, the complete observed data for subject $i$ is $(T_{i},\delta_{i},y_{i},t_{i})$. We denote by $m_{i}(t)$ the true and unobserved value of the longitudinal outcome at time $t$. To assess its effect on event risk, we estimate $m_{i}(t)$ via a mixed-effects model, which combines fixed effects to capture population-level trends and random effects to characterize individual variation, a well-established approach in longitudinal data analysis [laird1982random, verbeke2000linear]. Specifically, we incorporate quadratic splines of time to flexibly model subject-specific trajectories:
@@ 2.2.1 Joint modeling of longitudinal and survival data
Crucially, the hazard depends on the true unobserved longitudinal process $m_{i}(t)$ rather than the noisy observations $y_{i}(t)$. This connection is established through the random effects $b_{i}$, which serve as the bridge between the longitudinal and survival submodels, capturing how individual-specific biomarker trajectories relate to their risk of experiencing the event. For the baseline hazard function, several specifications are possible. In this study, $h_{0}(t)$ is modeled parametrically using the Weibull distribution [kalbfleisch2002statistical]. Alternatively, one could use other distributions or leave $h_{0}(t)$ unspecified, as in the Cox proportional hazards regression.
A classical approach to estimating joint model parameters relies on maximum likelihood, where the log-likelihood is derived from the joint distribution of the longitudinal and time-to-event outcomes. Obtaining the marginal likelihood requires approximation due to intractable integrals. Methods like Gaussian quadrature or Monte Carlo integration are commonly used [henderson2000joint, song2002semiparametric]. In this work, we instead adopt a Bayesian approach [rizopoulos2011dynamic, rizopoulos2012joint], where inference is based on the posterior distribution of the model parameters and latent variables. Specifically, we use Markov Chain Monte Carlo (MCMC) methods to estimate the parameters of the proposed joint model. The joint likelihood is formulated under the assumption that the longitudinal and survival processes are conditionally independent given the subject-specific random effects.
@@
2.2.2 Model specification
The longitudinal submodel, describing the evolution of log-transformed NT-proBNP values over time, was implemented via the nlme [pinheiro2017package] package. To flexibly capture the non-linear subject-specific trajectories and variability over time (see Figure 1F), we used quadratic splines of time for both fixed and random effects. We did not include additional covariates in the longitudinal model, in order to focus on the temporal dynamics of NT-proBNP and its association with event risk, and to avoid overfitting given the limited number of measurements for many patients. The survival submodel includes all pre-selected baseline covariates. We assumed a parametric Weibull baseline hazard to enable fully parametric estimation of the time-to-event distribution, implemented using the survival package [therneau2015package].
@@ 2.2.2 Model specification
For benchmarking purposes, we also fitted a standard survival Weibull model using all baseline covariates and the most recent NT-proBNP measurement as an additional covariate, consistent with the approach in prior work [zhang2018], and refer to this model as the benchmark model. This comparison is intended to determine whether modeling the entire biomarker trajectory dynamically offers advantages over relying on a single static measurement. All models were estimated using the JMbayes2[rizopoulos2025jmbayes2] package in R, which employs a full Bayesian framework with MCMC algorithms. Posterior inference is obtained via a combination of Gibbs sampling and Metropolis-Hastings algorithm. We used the default priors in the package JMbayes2, which primarily consist of weakly informative normal priors for the regression coefficients and half-t or gamma priors for the variance and standard deviation components.
3 Results
3.1 Exploratory data analysis
@@
The NT-proBNP measurements in this cohort span a wide range, with a median concentration of 1757 pg/mL (IQR: 578 - 4637.5 pg/mL). In our analysis NT-proBNP values are log-transformed to ensure positivity and to better approximate a symmetric distribution, following common practice in the literature [vanBoven2018, zhang2018]. Figure 1C displays the distribution of the log-transformed NT-proBNP measurements collected from any patient, totaling 9,594 values. The time distribution of these values is shown in Figure 1D, which illustrates that the time measurements of NT-proBNP are concentrated in the first part of the follow-up period, with almost 70% of the values recorded within the first 180 days after $t_{0}$. Recall that biomarker measurements taken within 30 days prior to $t_{0}$ and at $t_{0}$ are included in the analysis; in fact, approximately 31% of all measurements fall within this pre-baseline window. The average time of the first NT-proBNP measurement per patient is 9 days before $t_{0}$, while the average time of the last measurement is 367 days after $t_{0}$. The distribution of the number of NT-proBNP measurements per patient (Figure 1E) indicates that the largest proportion of patients (33%) have exactly three recorded values, which is the minimum required by our inclusion criteria. The limited number of longitudinal measurements represents a key challenge of the analysis, as it constrains our ability to robustly characterize individual biomarker trajectories over time. Individual longitudinal trajectories, shown in Figure 1F further demonstrate substantial heterogeneity in NT-proBNP patterns.
3.2 Model fit and baseline risk associations